# A matrix-free convex limiting framework for continuous Galerkin methods with nonlinear stabilization

### v1

@@ 1 Introduction
Finite element methods based on continuous Galerkin (CG) approximations are rarely used for discretization of nonlinear hyperbolic problems. It is commonly believed that finite volume and discontinuous Galerkin (DG) methods are better suited for this purpose because of the ease with which they can be manipulated to satisfy physical and numerical admissibility conditions. In particular, the use of slope limiters makes it possible to enforce preservation of local and/or global bounds for scalar quantities of interest [5, 24, 64]. Since slope limiting does not preserve global continuity, high-resolution finite element schemes of CG type are usually stabilized using nonlinear artificial viscosity [9, 10, 19, 21] or algebraic flux correction (AFC) tools [27, 30].
@@
Many classical AFC schemes for low-order (linear or multilinear) continuous finite elements belong to the family of flux-corrected transport (FCT) methods [31, 41]. In the scalar case, flux limiting can be performed using Zalesak’s algorithm [62] or a localized limiter for individual edge or element contributions [7, 12, 40]. Both approaches can be generalized to systems in a manner that guarantees the validity of positivity constraints for quasi-concave scalar functions of conserved variables [8, 9, 39]. Numerical schemes that provably satisfy such inequality constraints are called invariant domain preserving (IDP) [11] or, informally, positivity preserving [64].
@@
The localized FCT method proposed by Guermond et al. [9] constrains forward Euler stages of an explicit strong stability preserving Runge–Kutta (SSP-RK) method using convex limiting. This terminology reflects the fact that the nodal states of the flux-corrected CG approximation represent convex combinations of admissible states. The IDP property of the low-order (local Lax–Friedrichs) method, and of the FCT extension, is guaranteed under a CFL-like time step restriction. The monolithic convex limiting (MCL) procedure introduced in [25] can be combined with any time stepping method [35, 45, 50] and has been successfully extended to high-order finite elements [33, 34]. Convex limiting of FCT or MCL type is also possible for nodal DG methods using Bernstein or Legendre–Gauss–Lobatto (LGL) finite elements of arbitrary order [14, 15, 38, 47, 51]. However, achieving optimal accuracy and performance with such AFC schemes typically requires the use of compact-stencil representations, subcell flux limiters, and local bounds depending on ad hoc smoothness indicators (as shown already in [16, 17, 34, 40]).
@@
A powerful alternative framework for the design of positivity-preserving high-order DG methods was developed by Zhang and Shu [63, 64, 65]. In a typical implementation of their approach, which builds on the seminal paper [48] by Perthame and Shu, spurious oscillations are suppressed by using a weighted essentially nonoscillatory (WENO) reconstruction as a substitute for an unacceptable DG approximation in a troubled cell [43, 49, 52, 66, 67, 68, 69]. Positivity preservation can then be enforced using slope or flux limiting. The Zhang–Shu slope limiter ensures positivity at the quadrature points corresponding to a convex decomposition of the cell average into states that are evolved using a one-dimensional IDP scheme [27, 64]. Flux limiting makes it possible to constrain the cell averages directly [44, 60, 61]. In particular, FCT and MCL algorithms can be used for this purpose [13, 26, 27].
@@
The objective of this work is to develop a CG version of the Zhang–Shu limiting machinery for DG discretizations of hyperbolic problems. The first step toward that end has already been made in [36], where we designed nonlinear stabilization terms for CG discretizations using shock detectors that depend on deviations from a Hermite WENO reconstruction. In contrast to traditional WENO limiters, our approach is suitable for baseline discretizations of CG type and leads to analyzable nonlinear schemes [56]. It remains to design a limiter that ensures positivity preservation. The convex limiting strategy that we propose below is based on a representation of the CG approximation in terms of constrained nodal states. An element-based slope limiter controls deviations of these states from intermediate cell averages that are positivity preserving under a restrictive CFL condition. The optional use of a flux limiter enables us to guarantee positivity preservation under the mild CFL condition of the piecewise-constant DG-$\mathbb{P}_{0}$ approximation. Moreover, all calculations can be performed in a matrix-free manner, i.e., without calculating the full element matrices of the high-order finite element space. This desirable property of stabilized schemes paves the way for achieving high performance in hardware-oriented implementations [4, 6, 16, 17, 22].
@@
In Section 2, we discretize a generic hyperbolic problem in space using Bernstein finite elements and the CG-WENO method introduced in [36]. As we show in Section 3, the stabilized baseline scheme can be decomposed into a low-order IDP approximation and a sum of element contributions that may require limiting. Representing the low-order part in terms of intermediate cell averages, we formulate sufficient conditions for explicit SSP-RK time discretizations to be IDP. Monolithic limiters for auxiliary cell averages and auxiliary nodal states are presented in Section 4. Finally, we perform numerical studies and draw conclusions in Sections 5 and 6, respectively.
@@ 2 Stabilized CG discretization
where $u_{0}$ is the initial data, $T>0$ is the final time, and $\hat{u}$ is the data of a weakly imposed boundary condition. We denote by $\mathcal{F}(u_{L},u_{R};\mathbf{n})$ a boundary flux in the direction of the unit outward normal $\mathbf{n}$. The states $u_{L}$ and $u_{R}$ represent the initial states of the corresponding Riemann problem [27].
@@ 2.1 Galerkin scheme using Bernstein finite elements
To solve (1) using a CG-type finite element method, we use a conforming, possibly unstructured, mesh $\mathcal{T}_{h}$ consisting of simplices or $d$-dimensional boxes $K_{e},\ e=1,\ldots,E_{h}$. For simplicity, we assume that $\bar{\Omega}=\bigcup_{e=1}^{E_{h}}K_{e}$. A globally continuous approximation $u_{h}$ is sought in a space $V_{h}$ that is spanned by nodal basis functions $\varphi_{1},\ldots,\varphi_{N_{h}}\in C(\bar{\Omega})$ of polynomial degree $p\in\mathbb{N}$. The corresponding nodal points are denoted by $\mathbf{x}_{1},\ldots,\mathbf{x}_{N_{h}}$. The basis functions $\varphi_{i}$ of the Bernstein polynomial basis are nonnegative and form a partition of unity, i.e., $\sum_{i=1}^{N_{h}}\varphi_{i}\equiv 1$. For a formal definition of $\varphi_{i}$ and a review of further properties, we refer the reader to [14, 33, 34, 40] and [27, Sec. 6.1].
@@ 2.2 Dissipative WENO stabilization
on the left-hand side of (5). The CG-WENO methods presented in [36, 56] blend the local bilinear forms of high- and low-order stabilization operators using a smoothness sensor $\gamma_{e}\in[0,1]$. The element contributions
@@
are defined using the consistent $L^{2}$ projection $\mathbf{g}_{h}$ of $\nabla u_{h}$ into the CG space. The viscosity parameter $\nu_{e}=\frac{\lambda_{e}h_{e}}{2p}$ depends on the local mesh size $h_{e}$, polynomial degree $p$, and a local bound $\lambda_{e}$ for the maximum speed. Details regarding the calculation of $\mathbf{g}_{h}$ and $\lambda_{e}$ can be found in [33, 36, 40, 56].
In formula (6), the smoothness sensor $\gamma_{e}$ acts as a slope limiter for the projected gradient $\mathbf{g}_{h}$. In the case $\gamma_{e}=0$, the local bilinear form (6) introduces low-order stabilization of local Lax–Friedrichs type. The setting $\gamma_{e}=1$ corresponds to the linear high-order stabilization employed in [33, 40]. The symmetry and positive semi-definiteness of the corresponding bilinear form can be shown as in [46, Lem. 1]. The WENO shock detector
@@
proposed in [36] measures deviations of $u_{h}^{e}=u_{h}|_{K_{e}}$ from a WENO reconstruction $u_{h}^{e,*}$. The sensitivity of $\gamma_{e}$ to these deviations can be varied by adjusting the parameter $q\geq 1$. The Sobolev semi-norm $\|\cdot\|_{e}$ is defined by
@@
Following the design of Hermite WENO limiters for DG methods [43, 67, 69], the reconstruction $u_{h}^{e,*}$ is defined as a convex combination of candidate polynomials $u_{h,l}^{e},\ l=0,\ldots,n_{e}$ with nonlinear weights [20]
@@
that depend on the choice of the linear weights $w_{l}^{e,\rm lin}$ and on the power $r\in\mathbb{N}$. The small parameter $\epsilon>0$ is used to avoid division by zero. The candidate polynomials $u_{h,l}^{e}$ are constructed using Hermite interpolation from cells $K_{e^{\prime}}$ sharing a facet $S_{ee^{\prime}}$ with $K_{e}$. For details, we refer the reader to [36, 56].
@@ 3 Low-order IDP discretization
In many cases, the results produced by the fully discrete version of the CG-WENO scheme (9) are excellent [36, 56]. However, there is no guarantee that the nodal states $u_{j}$ of the Bernstein finite element approximation are positivity preserving / IDP. To show that the IDP property can be enforced using monolithic convex limiting, we first approximate $\mathbf{f}(u_{h})$ by
@@
of the semi-discrete equation (10) using (cf. [27, Ch. 3 and 6], [29])
@@
$b_{i}(u_{h},\hat{u}_{h})$ $=\sum_{e\in\mathcal{E}_{i}}\sum_{e^{\prime}\in\mathcal{B}_{e}^{\partial}}\sigma_{i,ee^{\prime}}[\mathcal{F}(u_{i},\hat{u}_{i,ee^{\prime}};\mathbf{n})-\mathbf{f}(u_{i})],$ (16)
@@
The stabilization term $s_{i}(u_{h})$ introduces low-order Rusanov dissipation (cf. [1], [27, Ch. 4]), the levels of which are inversely proportional to a ‘fake’ time step $\Delta t_{e}$. An IDP upper bound for $\Delta t_{e}$ is derived below. The nonlinear term $a_{i}(u_{h})$ of the evolution equation (12) approximates $a(\varphi_{i},u_{h})$ using (11) and inexact quadrature for the integral $\int_{K_{e}}\varphi_{i}\nabla\cdot\mathbf{f}_{h}\mathrm{d}\mathbf{x}$. The approximation of $(\varphi_{i},u_{h})$ by $m_{i}u_{i}$ corresponds to mass lumping, which is equivalent to using inexact quadrature for $L^{2}$ scalar products. For Bernstein finite elements $m_{i}^{e}=\frac{|K_{e}|}{|\mathcal{N}_{e}|}$, where $|\mathcal{N}_{e}|$ is the cardinality of the index set $\mathcal{N}_{e}$ and $|K_{e}|$ is the $d$-dimensional volume of $K_{e}$. Similar mass lumping is performed in the boundary term (16), in which we use the local Lax–Friedrichs (LLF) flux
@@ 3.1 Intermediate cell averages
where $\Delta t_{e}$ is the fake time step of the Rusanov dissipation (15) and
@@
Under this condition, (22) yields a convex combination $u_{i}^{\rm SSP}$ of the states $u_{i}$ and $\bar{u}^{e},\ e\in\mathcal{E}_{i}$. If these states belong to a convex admissible set $\mathcal{G}$, then so does $u_{i}^{\rm SSP}$. Hence, the choice of $\Delta t_{e}$ should ideally guarantee that
$u_{j}\in\mathcal{G}\quad\forall j\in\mathcal{N}_{e}\quad\Rightarrow\quad\bar{u}^{e}\in\mathcal{G}.$ (24)
For an element $K_{e}=[x_{i},x_{j}],\ j=i+1$ of a 1D mesh with local spacing $h_{e}=|K_{e}|$, the intermediate state $\bar{u}^{e}$ defined by (20) reduces to
$\bar{u}_{ij}=\frac{u_{i}+u_{j}}{2}-\frac{\Delta t_{e}}{h_{e}}[f(u_{j})-f(u_{i})].$ (25)
Representation of (semi-)discrete schemes in terms of such bar states provides a convenient framework for proving the IDP property of first-order schemes and convex limiting algorithms [9, 11, 18, 25, 27]. Suppose that
$\lambda_{ij}\frac{\Delta t_{e}}{h_{e}}\leq\frac{1}{2},$ (26)
where $\lambda_{ij}$ is a local bound for the maximum speed. Then the interpretation of $\bar{u}_{ij}$ as a spatially averaged exact solution to the Riemann problem with initial states $u_{i}$ and $u_{j}$ proves that $\bar{u}_{ij}\in\mathcal{G}$ whenever $u_{i},u_{j}\in\mathcal{G}$.
In the multidimensional case, the bar states $\bar{u}_{ij}$ and their flux-corrected counterparts are associated with pairs of neighbor nodes or cells. The representation of the SSP-RK stage (22) in terms of element-based intermediate states $\bar{u}^{e}$ rather than their edge-based counterparts $\bar{u}_{ij}$ opens a new avenue for analysis and design of multidimensional IDP schemes. Remarkably, formula (20) resembles an explicit update for a cell average of a DG approximation. Taking advantage of this fact, we will adapt the analysis of positivity-preserving DG schemes [63, 64, 65] to our continuous finite element approximations.
3.2 Conditional IDP properties
It turns out that the IDP condition (24) is, indeed, satisfied if $\Delta t_{e}$ is sufficiently small. We state and prove this result in the following lemma.

### v2

# An element-based convex limiting framework for continuous Galerkin methods with nonlinear stabilization

@@ 1 Introduction
Finite element methods based on continuous Galerkin (CG) approximations are rarely used for discretization of nonlinear hyperbolic problems. It is commonly believed that finite volume and discontinuous Galerkin (DG) methods are better suited for this purpose because of the ease with which they can be manipulated to satisfy physical and numerical admissibility conditions. In particular, the use of slope limiters makes it possible to enforce preservation of local and/or global bounds for scalar quantities of interest [barth1989, kuzmin2010, zhang2011]. Since slope limiting does not preserve global continuity, high-resolution finite element schemes of CG type are usually stabilized using nonlinear artificial viscosity [guermond2018, guermond2011, hughes2010, johnson1990] or algebraic flux correction (AFC) tools [kuzmin2023, kuzmin2012b].
@@
Many classical AFC schemes for low-order (linear or multilinear) continuous finite elements belong to the family of flux-corrected transport (FCT) methods [kuzmin2010a, lohner1987]. In the scalar case, flux limiting can be performed using Zalesak’s algorithm [zalesak1979] or a localized limiter for individual edge or element contributions [cotter2016, guermond2017, lohmann2017]. Both approaches can be generalized to systems in a manner that guarantees the validity of positivity constraints for quasi-concave scalar functions of conserved variables [dobrev2018, guermond2018, lohmann2016]. Numerical schemes that provably satisfy such inequality constraints are called invariant domain preserving (IDP) [guermond2016] or, informally, positivity preserving [zhang2011].
@@
The localized FCT method proposed by Guermond et al. [guermond2018] constrains forward Euler stages of an explicit strong stability preserving Runge–Kutta (SSP-RK) method using convex limiting. This terminology reflects the fact that the nodal states of the flux-corrected CG approximation represent convex combinations of admissible states. The IDP property of the low-order (local Lax–Friedrichs) method, and of the FCT extension, is guaranteed under a CFL-like time step restriction. The monolithic convex limiting (MCL) procedure introduced in [kuzmin2020] can be combined with any time stepping method [kuzmin2022a, moujaes2025, quezada-de-luna2022] and has been successfully extended to high-order finite elements [kuzmin2020f, kuzmin2020a]. Convex limiting of FCT or MCL type is also possible for nodal DG methods using Bernstein or Legendre–Gauss–Lobatto (LGL) finite elements of arbitrary order [hajduk2021, hajduk2025a, lin2024, pazner2021, rueda-ramirez2024]. However, achieving optimal accuracy and performance with such AFC schemes typically requires the use of compact-stencil representations, subcell flux limiters, and local bounds depending on ad hoc smoothness indicators (as shown already in [hajduk2020, hajduk2020b, kuzmin2020a, lohmann2017]).
@@
A powerful alternative framework for the design of positivity-preserving high-order DG methods was developed by Zhang and Shu [zhang2010b, zhang2011, zhang2012]. In a typical implementation of their approach, which builds on the seminal paper [perthame1996] by Perthame and Shu, spurious oscillations are suppressed by using a weighted essentially nonoscillatory (WENO) reconstruction as a substitute for an unacceptable DG approximation in a troubled cell [luo2007, qiu2005, shu2016, zhong2013, zhu2009, zhu2008, zhu2017]. Positivity preservation can then be enforced using slope or flux limiting. The Zhang–Shu slope limiter ensures positivity at the quadrature points corresponding to a convex decomposition of the cell average into states that are evolved using a one-dimensional IDP scheme [kuzmin2023, zhang2011]. Flux limiting makes it possible to constrain the cell averages directly [moe2017, xiong2016, xu2014]. In particular, FCT and MCL algorithms can be used for this purpose [guermond2019, kuzmin2021, kuzmin2023].
@@
The objective of this work is to show that the methodology developed by Zhang and Shu can be adapted to continuous finite element approximations. The first step toward that end has already been made in [kuzmin2023a], where we designed nonlinear stabilization terms for CG discretizations using shock detectors that depend on deviations from a Hermite WENO reconstruction. In contrast to traditional WENO limiters, our approach is suitable for baseline discretizations of CG type and leads to analyzable nonlinear schemes [vedral2025]. It remains to design a limiter that ensures positivity preservation. The convex limiting strategy that we propose below is based on a representation of the CG approximation in terms of constrained nodal states. An element-based slope limiter controls deviations of these states from intermediate cell averages that are positivity preserving under a restrictive CFL condition. The optional use of a flux limiter enables us to guarantee positivity preservation under the mild CFL condition of the piecewise-constant DG-$\mathbb{P}_{0}$ approximation. Moreover, all calculations can be performed in a matrix-free manner, i.e., without calculating the full element matrices of the high-order finite element space.
@@
In Section 2, we discretize a generic hyperbolic problem in space using Bernstein finite elements and the CG-WENO method introduced in [kuzmin2023a]. As we show in Section 3, the stabilized baseline scheme can be decomposed into a low-order IDP approximation and a sum of element contributions that may require limiting. Representing the low-order part in terms of intermediate cell averages, we formulate sufficient conditions for explicit SSP-RK time discretizations to be IDP. Monolithic limiters for auxiliary cell averages and auxiliary nodal states are presented in Section 4. Finally, we perform numerical studies and draw conclusions in Sections 5 and 6, respectively.
@@ 2 Stabilized CG discretization
where $u_{0}$ is the initial data, $T>0$ is the final time, and $\hat{u}$ is the data of a weakly imposed boundary condition. We denote by $\mathcal{F}(u_{L},u_{R};\mathbf{n})$ a boundary flux in the direction of the unit outward normal $\mathbf{n}$. The states $u_{L}$ and $u_{R}$ represent the initial states of the corresponding Riemann problem [kuzmin2023].
@@ 2.1 Galerkin scheme using Bernstein finite elements
To solve (1) using a CG-type finite element method, we use a conforming, possibly unstructured, mesh $\mathcal{T}_{h}$ consisting of simplices or $d$-dimensional boxes $K_{e},\ e=1,\ldots,E_{h}$. For simplicity, we assume that $\bar{\Omega}=\bigcup_{e=1}^{E_{h}}K_{e}$. A globally continuous approximation $u_{h}$ is sought in a space $V_{h}$ that is spanned by nodal basis functions $\varphi_{1},\ldots,\varphi_{N_{h}}\in C(\bar{\Omega})$ of polynomial degree $p\in\mathbb{N}$. The corresponding nodal points are denoted by $\mathbf{x}_{1},\ldots,\mathbf{x}_{N_{h}}$. The basis functions $\varphi_{i}$ of the Bernstein polynomial basis are nonnegative and form a partition of unity, i.e., $\sum_{i=1}^{N_{h}}\varphi_{i}\equiv 1$. For a formal definition of $\varphi_{i}$ and a review of further properties, we refer the reader to [hajduk2021, kuzmin2020f, kuzmin2020a, lohmann2017] and [kuzmin2023, Sec. 6.1].
@@
$e\in\{1,\ldots,E_{h}\}$ index of a mesh element $K_{e}$ belonging to $\mathcal{T}_{h}$
$e^{\prime}\in\mathcal{B}_{e}$ index of a facet { $\begin{cases}\mbox{ $S_{ee^{\prime}}=\partial K_{e}\cap\partial K_{e^{\prime}}$ if $e^{\prime}\leq E_{h}$}\\ \mbox{ $S_{ee^{\prime}}\subset\partial K_{e}\cap\partial\Omega$ if $e^{\prime}>E_{h}$}\end{cases}$
$e^{\prime}\in\mathcal{B}_{e}^{\partial}$ index of a boundary facet $S_{ee^{\prime}}$ with $e^{\prime}>E_{h}$
$i\in\{1,\ldots,N_{h}\}$ index of a nodal point $\mathbf{x}_{i}$ or basis function $\varphi_{i}$
$i\in\mathcal{N}_{e}$ index of a nodal point $\mathbf{x}_{i}\in K_{e}$
$i\in\mathcal{N}_{e}^{\partial}$ index of a boundary node $\mathbf{x}_{i}\in\partial K_{e}$
[…]
$e\in\mathcal{E}_{i}$ index of a mesh element $K_{e}$ containing $\mathbf{x}_{i}$
$j\in\mathcal{J}_{i}$ index of a point $\mathbf{x}_{j}\in K_{e}$ for some $e\in\mathcal{E}_{i}$
Table 1: Index notation used in descriptions of numerical schemes.
@@ 2.2 Dissipative WENO stabilization
on the left-hand side of (5). The CG-WENO methods presented in [kuzmin2023a, vedral2025] blend the local bilinear forms of high- and low-order stabilization operators using a smoothness sensor $\gamma_{e}\in[0,1]$. The element contributions
@@
are defined using the consistent $L^{2}$ projection $\mathbf{g}_{h}$ of $\nabla u_{h}$ into the CG space. The viscosity parameter $\nu_{e}=\frac{\lambda_{e}h_{e}}{2p}$ depends on the local mesh size $h_{e}$, polynomial degree $p$, and a local bound $\lambda_{e}$ for the maximum speed. Details regarding the calculation of $\mathbf{g}_{h}$ and $\lambda_{e}$ can be found in [kuzmin2020f, kuzmin2023a, lohmann2017, vedral2025].
In formula (6), the smoothness sensor $\gamma_{e}$ acts as a slope limiter for the projected gradient $\mathbf{g}_{h}$. In the case $\gamma_{e}=0$, the local bilinear form (6) introduces low-order stabilization of local Lax–Friedrichs type. The setting $\gamma_{e}=1$ corresponds to the linear high-order stabilization employed in [kuzmin2020f, lohmann2017]. The symmetry and positive semi-definiteness of the corresponding bilinear form can be shown as in [olshanskii2025, Lem. 4.1]. The WENO shock detector
@@
proposed in [kuzmin2023a] measures deviations of $u_{h}^{e}=u_{h}|_{K_{e}}$ from a WENO reconstruction $u_{h}^{e,*}$. The sensitivity of $\gamma_{e}$ to these deviations can be varied by adjusting the parameter $q\geq 1$. As $q$ is increased, the resolution of steep gradients improves, but violations of maximum principles may occur / become more pronounced (see Section 5.3). The Sobolev semi-norm $\|\cdot\|_{e}$ is defined by
@@
Following the design of Hermite WENO limiters for DG methods [luo2007, zhu2009, zhu2017], the reconstruction $u_{h}^{e,*}$ is defined as a convex combination of candidate polynomials $u_{h,l}^{e},\ l=0,\ldots,n_{e}$ with nonlinear weights [jiang1996]
@@
that depend on the choice of the linear weights $w_{l}^{e,\rm lin}$. The small parameter $\epsilon>0$ is used to avoid division by zero. The exponent $r\geq 1$ determines the sensitivity of the nonlinear weights to the smoothness indicators. Larger values of the parameter $r$ increase the difference between the weights of smooth and nonsmooth candidates but can degrade the accuracy in smooth regions. Liu et al. [liu1994] suggest setting $r$ equal to the polynomial degree of the WENO reconstruction. Jiang and Shu [jiang1996] found that $r=2$ achieves the best balance between robustness and accuracy. This is currently the default choice in classical WENO schemes. In our implementation, the candidate polynomials $u_{h,l}^{e}$ are constructed using Hermite interpolation from cells $K_{e^{\prime}}$ sharing a facet $S_{ee^{\prime}}$ with $K_{e}$. For details, we refer the reader to [kuzmin2023a, vedral2025].
The design of the local stabilization term (6) is motivated by the fact that deviations of $u_{h}^{e}$ from $u_{h}^{e,*}$ are large in regions where $u_{h}$ is not smooth. In these regions, the value of the WENO sensor $\gamma_{e}$ is small, and (6) introduces large amounts of low-order numerical viscosity. In smooth regions, the value of $\gamma_{e}$ is close to unity. The high-order stabilization corresponding to $\gamma_{e}=1$ ensures optimal convergence to smooth exact solutions [kuzmin2020f, lohmann2017].
@@ 3 Low-order IDP discretization
In many cases, the results produced by the fully discrete version of the CG-WENO scheme (9) are excellent [kuzmin2023a, vedral2025]. However, there is no guarantee that the nodal states $u_{j}$ of the Bernstein finite element approximation are positivity preserving / IDP. To show that the IDP property can be enforced using monolithic convex limiting, we first approximate $\mathbf{f}(u_{h})$ by
@@
of the semi-discrete equation (10) using (cf. [kuzmin2023, Ch. 3 and 6], [kuzmin2025a])
@@
$b_{i}(u_{h},\hat{u}_{h})$ $=\sum_{e\in\mathcal{E}_{i}}\sum_{e^{\prime}\in\mathcal{B}_{e}^{\partial}}\sigma_{i,ee^{\prime}}[\mathcal{F}(u_{i},\hat{u}_{i,ee^{\prime}};\mathbf{n})-\mathbf{f}(u_{i})\cdot\mathbf{n}],$ (16)
@@
The stabilization term $s_{i}(u_{h})$ introduces low-order Rusanov dissipation (cf. [abgrall2006], [kuzmin2023, Ch. 4]), the levels of which are inversely proportional to a ‘fake’ time step $\Delta t_{e}$. An IDP upper bound for $\Delta t_{e}$ is derived below. The nonlinear term $a_{i}(u_{h})$ of the evolution equation (12) approximates $a(\varphi_{i},u_{h})$ using (11) and inexact quadrature for the integral $\int_{K_{e}}\varphi_{i}\nabla\cdot\mathbf{f}_{h}\mathrm{d}\mathbf{x}$. The approximation of $(\varphi_{i},u_{h})$ by $m_{i}u_{i}$ corresponds to mass lumping, which is equivalent to using inexact quadrature for $L^{2}$ scalar products. For Bernstein finite elements $m_{i}^{e}=\frac{|K_{e}|}{|\mathcal{N}_{e}|}$, where $|\mathcal{N}_{e}|$ is the cardinality of the index set $\mathcal{N}_{e}$ and $|K_{e}|$ is the $d$-dimensional volume of $K_{e}$. Similar mass lumping is performed in the boundary term (16), in which we use the local Lax–Friedrichs (LLF) flux
@@ 3.1 Intermediate cell averages
where $\Delta t_{e}$ is the pseudo time step of the Rusanov dissipation (15) and
@@
Under this condition, (22) yields a convex combination $u_{i}^{\rm SSP}$ of the states $u_{i}$ and $\bar{u}^{e},\ e\in\mathcal{E}_{i}$. If these states belong to a convex admissible set $\mathcal{G}$, then so does $u_{i}^{\rm SSP}$ and, therefore, the SSP-RK update is IDP w.r.t. $\mathcal{G}$.
3.2 Analysis in the one-dimensional case
For an element $K_{e}$ of a 1D mesh with $\mathcal{N}_{e}=\{i,j\}$ and local spacing $h_{e}=|K_{e}|$, the intermediate state $\bar{u}^{e}$ defined by (20) reduces to
$\bar{u}_{ij}=\frac{u_{i}+u_{j}}{2}-(j-i)\frac{\Delta t_{e}}{h_{e}}[f(u_{j})-f(u_{i})]=\bar{u}_{ji}.$ (24)
Let $\lambda_{ij}^{\max}$ denote the maximum speed of the 1D Riemann problem with the initial states $u_{i}$ and $u_{j}$. If the pseudo time step $\Delta t_{e}$ satisfies
$\frac{\lambda_{ij}^{\max}\Delta t_{e}}{h_{e}}\leq\frac{1}{2},$ (25)
then the bar state $\bar{u}_{ij}$ defined by (24) corresponds to a spatially averaged exact solution to the Riemann problem [guermond2016, harten1983a]. Thus, $u_{i},u_{j}\in\mathcal{G}$ implies $\bar{u}_{ij}\in\mathcal{G}$ for every invariant domain $\mathcal{G}$ of the hyperbolic problem.
The 1D version of (22) can be rewritten in terms of $\bar{u}_{ij}$ as follows:
$u_{i}^{\rm SSP}=\Big(1-\frac{\Delta t}{m_{i}}\sum_{j\in\mathcal{J}_{i}\backslash\{i\}}2d_{ij}\Big)u_{i}+\frac{\Delta t}{m_{i}}\sum_{j\in\mathcal{J}_{i}\backslash\{i\}}2d_{ij}\bar{u}_{ij}.$ (26)
Here $\mathcal{J}_{i}=\mathcal{N}_{e^{\prime}}\cup\mathcal{N}_{e}$ for $\mathcal{N}_{e^{\prime}}=\{i-1,i\}$ and $\mathcal{N}_{e}=\{i,i+1\}$. The auxiliary coefficients $d_{i,i-1}=m_{i}^{e^{\prime}}/(2\Delta t_{e^{\prime}})$ and $d_{i,i+1}=m_{i}^{e}/(2\Delta t_{e})$ are positive.